OpenStudy (anonymous):

he value of 1/3*5+1/5*7+1/7*9+....+1/23*25 is

OpenStudy (tkhunny):

It may help to rewrite the terms a bit: Numbering the terms: 1: 1/(3*5) 2: 1/(5*7) 3: 1/(7*9) etc We can refer to the terms by their sequence numbers \(\dfrac{1}{(2(1)+1)\cdot (2(1) + 3)} + \dfrac{1}{(2(2)+1)\cdot (2(2) + 3)} + ... etc\) Or generally, we have for n = 1, 2, 3, ... 11, just \(\dfrac{1}{(2n+1)\cdot (2n + 3)}\) That doesn't do much until you find that: \(\dfrac{1}{(2n+1)\cdot (2n + 3)} = \dfrac{1/2}{2n+1} - \dfrac{1/2}{2n+3}\). If you write all your terms like that last expression, something wonderful will occur.